# Finer: Investigating and Enhancing Fine-Grained Visual Concept Recognition in Large Vision Language Models

### v2

@@ Abstract
Recent advances in instruction-tuned Large Vision-Language Models (LVLMs) have imbued the models with the ability to generate high-level, image-grounded explanations with ease. While such capability is largely attributed to the rich world knowledge contained within the Large Language Models (LLMs), our work reveals their shortcomings in fine-grained visual categorization (FGVC) across six different benchmark settings. Most recent state-of-the-art LVLMs like LLaVa-1.5, InstructBLIP and GPT-4V not only severely deteriorate in terms of classification performance, e.g., average drop of 65.58 in EM for Stanford Dogs for LLaVA-1.5, but also struggle to generate an accurate explanation with detailed attributes based on the concept that appears within an input image despite their capability to generate holistic image-level descriptions. In-depth analyses show that instruction-tuned LVLMs exhibit modality gap, showing discrepancy when given textual and visual inputs that correspond to the same concept, preventing the image modality from leveraging the rich parametric knowledge within the LLMs. In an effort to further the community’s endeavor in this direction, we propose a multiple granularity attribute-centric evaluation benchmark, Finer, which aims to establish a ground to evaluate LVLMs’ fine-grained visual comprehension ability and provide significantly improved explainability. [^1] We release the dataset and code at: https://github.com/wjdghks950/Finer
@@ 1 Introduction
The surge of interest and development of Large Vision-Language Models (LVLMs) that are able to generate image-grounded texts have seen significant progress. Models such as InstructBLIP (Dai et al., 2023) and LLaVA (Liu et al., 2023b, a) have consistently exhibited strong zero-shot capability in generating image captions, visual reasoning and textual descriptions, and even leveraging external knowledge for complex question answering tasks Marino et al. (2019); Schwenk et al. (2022). Such results across diverse benchmarks indicate that these models, most of them being built on large language models (LLMs) like Vicuna (Chiang et al., 2023), Flan-T5 (Chung et al., 2022), Llama (Touvron et al., 2023), are already equipped with the ability to simultaneously leverage the interplay between the textual knowledge acquired during pre-training and the image understanding ability acquired during instruction-tuning. Furthermore, all of these models exhibit strong zero-shot task transferability to multiple downstream tasks.
@@
Conventionally, in the image classification field, many previous works on fine-grained visual classification (FGVC) (Wei et al., 2022b; Diao et al., 2022; Zhu et al., 2022; Yang et al., 2022; Wang et al., 2022) sought to accurately classify diverse images ranging from different types of birds, plants and animals Van Horn et al. (2015, 2018) and artificial objects such as cars Krause et al. (2013) and aircrafts Maji et al. (2013). In this work, we investigate whether state-of-the-art LVLMs can combine their image understanding ability and rich knowledge acquired from the Web to handle FGVC tasks. To our surprise, while the LVLMs perform almost perfectly, e.g., 98.43 for LLaVA-1.5 (13B) on iNaturalist, at the superordinate-level granularity (e.g., birds, jets), their classification abilities do not extend to the coarse and finer-grained concepts (e.g., bald eagle, F-22 Raptor), exhibiting substantially deteriorated classification performance (§3.2); 46.91 for coarse-level and 1.56 for fine-level categories on iNaturalist. Our empirical analyses of these models reveal that these seemingly strong models suffer from modality gap (Liang et al., 2022). We empirically demonstrate that such discrepancy stems from LVLMs projecting visual and textual modalities disparately, limiting them from leveraging rich parametric knowledge, obtained during the textual pre-training stage of LLMs, to infer fine-grained concepts. Our results also show that such constraints lead to diminished fine-grained comprehension of the image, preventing these models from generating high-quality, finer granularity description of the concepts that appear within an image.
@@
We also present an attribute-centric and multiple granularity image classification benchmark, Finer. Our benchmark broadens the scope of six widely-used FGVC benchmarks like iNaturalist Van Horn et al. (2018) and FGVC-Aircrafts Maji et al. (2013) by (i) generating multiple granularity levels for fine-grained visual concept recognition, and (ii) constructing a visual, concept-indicative attribute sets per fine-grained concept to measure the ability of LVLMs to generate fine-grained concept descriptions given an image. To summarize, our contributions include:
@@
  • We highlight the lack of fine-grained image comprehension ability of instruction-tuned LVLMs across various real-life objects. To the best of our knowledge, we are the first to explore FGVC as an evaluation criteria for these models.
@@
  • We highlight the persistence of modality gap within the LVLMs by conducting an extensive per-modality-based probing, revealing the discrepancy in how the two modalities are processed by these models.
@@
  • We construct a novel attribute-centric, multi-granularity benchmark on FGVC datasets to open up a new direction for future works to measure LVLMs’ modality gap and their fine-grained image understanding capability. Fine-tuning on Finer proves beneficial by closing the modality gap and improving FGVC accuracy for held-out datasets.
@@ 2 Related Work
2.1 Instruction-tuning of Large Vision- Language Models
The recent advances in LVLMs such as LLaVA (Liu et al., 2023a, b), BLIP-2 Li et al. (2023), InstructBLIP (Dai et al., 2023), and closed-source models like GPT-4V OpenAI (2023); Yang et al. (2023) have brought to our attention their zero-shot task solving abilities, especially in downstream benchmarks such as Visual Question Answering (VQA), reasoning and image captioning, all of which require output generation conditioned on the extensive knowledge of the real-world. Their reasoning and captioning abilities, based on an intricate interplay between their parametric knowledge and observation of images, result in sensible outputs. However, none of them evaluates their visual comprehension abilities to identify fine-grained objects that appear within an input image.
2.2 Vision-Language Alignment Models
Alignment-based VLMs like CLIP Radford et al. (2021) and GLIP (Li et al., 2022) rely on minimizing the contrastive objective between the visual and textual representations. A recent analytical work (Liang et al., 2022) on CLIP-like models reveals that there is a modality gap between the text and visual modalities, which imposes substantial implications on downstream task performances. Our work shows that such modality gap persists even after the image representations have been projected onto the textual representational space, which we hypothesize in this work as the underlying reason behind the substantial degradation in FGVC.
@@ 2.2 Vision-Language Alignment Models
2.3 Fine-grained Visual Categorization
Previous works approach FGVC with masked image modeling (He et al., 2022; Ryali et al., 2023) or incorporating concept-related meta-information (e.g., location, spatio-temporal prior, text description) (Diao et al., 2022) or LLM-generated textual descriptions of concepts (Menon and Vondrick, 2023). However, learning of fine-grained visual categories in LVLMs and their ability to elaborate on the fine-grained details of the input image via text generation is yet to be explored. Furthermore, recent works in FGVC have also shown to disregard fine-grained details of images (Krojer et al., 2022) and videos Wang et al. (2023), work poorly on downstream tasks involving localization (Ranasinghe et al., 2023; Zhong et al., 2022), object attributes (Yuksekgonul et al., 2022) and actions Wang et al. (2023).
Figure 3: Fine-grained classification pipeline. At each granularity level, a predicted output from LVLM is injected into the next level prompt. (1) Superordinate-level prompt is used to predict the highest-level category (e.g., bird). (2) Coarse-level prompt is subsequently fed with the predicted output and fed back to the LVLM to generate the next output (e.g., parrot), and (3a) and (4) follow the same steps. (3b) illustrates AttrSeek, an alternative path for (3a), wherein the model generates the attributes first.
@@ 3.1 Evaluation Settings
We assess the accuracy of the generated concept labels given an image and granularity-specific prompt asking the model to figure out the correct category the concept in the image belongs to. Following previous works on concept classification using auto-regressive models, we employ F1 and Exact Match (EM) scores; note that the EM score used in this work is a modified EM score that parses a sentence and considers the output label correct if the ground-truth label string exists within a pre-defined maximum number of tokens, $m$ (in this work, we set this value to $m=20$).
@@
The models used in this work are as follows: LLaVA-1.5 (Liu et al., 2023b, a), InstructBLIP (Dai et al., 2023), BLIP-2 (Li et al., 2023), and GPT-4V; the hyperparameter setting for each model is in Appendix A.1. The open-sourced models like LLaVA-1.5 and InstructBLIP follow a generic pipeline of transforming an input image $\mathbf{X_{v}}$ with a frozen vision encoder such as CLIP ViT-L/14 Radford et al. (2021) into an encoded image representation $\mathbf{Z_{v}}$. Then, these models either project $\mathbf{Z_{v}}$ into the language representation space through a learned projection layer $\mathbf{W}$, which becomes $\mathbf{H_{v}}=\mathbf{W}\cdot\mathbf{Z_{v}}$ as in Liu et al. (2023b), or attend over $\mathbf{Z_{v}}$ with learnable queries $\mathbf{Q}$ as in Li et al. (2023); Dai et al. (2023). Such transformed visual representations interact with $\mathbf{X_{instruct}}$, a language instruction, which attends over the image representations (or queries) within the self-attention layers of the LLM component, with the model generating the final output sequence.
@@ 3.2 Brittleness of Vision-Language Models
To evaluate the fine-grained image categorization ability of LVLMs, we measure their classification performance per granularity by prompting the models to generate the correct label for a given concept image as shown in Figure 2 . As elaborated in Figure 3, we assess the models’ classification ability across three different granularity levels. In Figure 2, we evidence significant deterioration in terms of classification performance across all the five baselines, with some even reaching near 0% in EM score. While models do perform very well for superordinate-level categories, with some reaching 100% in EM, the finer granularity leads to worse classification performance. In terms of model size, larger models like LLaVA-1.5 (13B) and GPT-4V tend to perform better than smaller models like the 7B variants. For InstructBLIP, the 7B version performs better than the 13B version by a large margin, with the 13B version exhibiting less capable instruction-following ability than the 7B one, potentially due to 7B variant being less prone to overfitting and exhibiting efficiency in simple tasks like classification.
@@
Table 1: Elicitive prompting results (%) on iNaturalist. Prompting techniques like Chain-of-Thought (CoT) on the fine-grained classification (Fine) is unable to improve performance in open-source models.
@@
LLMs like Vicuna (Chiang et al., 2023) and LLama (Touvron et al., 2023) used as textual reasoning components of VLMs are known to perform better when presented with elicitive prompts like Chain-of-Thought (CoT) (Wei et al., 2022a) that improve model’s reasoning ability. A unifying thought along this line of prompting techniques is to break down a complex problem into a sequence of sub-problems, i.e., divide-and-conquer. Inspired by these prompting techniques, we propose and evaluate our prompting technique for FGVC, AttrSeek. In this simple prompting strategy, we first (i) prompt the VLMs to generate the most distinctive physical attributes visible in the concepts in an image, and (ii) feed the generated set of attributes along with the concept-asking prompt for final prediction as shown in Figure 3. However, in Table 1, we do not see much improvement in terms of fine-grained concept classification in the open-source models, with neither CoT nor AttrSeek enhancing the classification performance. However, there is a substantial increase in fine-grained performance for GPT-4V when prompted with first with the attribute-seeking prompt and then predicting the end concept. This result suggests that LLaVA-1.5 and InstructBLIP, while they exhibit strong image captioning and reasoning ability, is limited in terms of image grounded attribute understanding even when provided with elicitive prompts. This result also suggests that open-source VLMs may lag behind in instruction-following abilities in comparison to GPT-4V For additional details on few-shot prompting, see Appendix A.3.
@@ 4 Modality Gap: Discrepancy Between Textual and Visual Modalities
We first hypothesize that the lack of zero-shot concept classification ability of LVLMs is mainly due to the modality gap of textual and visual inputs, preventing the models from leveraging the existing concept-related parametric knowledge when an image of a concept is given. In this section, we aim to delve into how these models process visual and textual modalities by exploring how well they perform when given only textual descriptions of a concept (§4.1) and how they accurately elaborate what they see in a given image (§4.2).
Base Model LLM Input Modality Similarity Measurement Metrics ROUGE-1 ROUGE-2 ROUGE-L BertScore AlignScore LLaVA-1.5 (Liu et al., 2023a) Vicuna-7B Text 22.343 11.023 21.161 85.429 18.935 Image 17.468 10.049 17.314 84.499 15.509 Vicuna-13B Text 20.616 10.902 19.647 85.944 15.746 Image 19.661 11.334 19.114 85.704 13.319 $\ \ \Delta$ Avg. 2.920 0.703 2.189 0.585 2.926 InstructBLIP (Dai et al., 2023) Vicuna-7B Text 18.987 9.192 16.577 83.102 15.002 Image 16.018 9.788 11.258 82.681 10.186 Vicuna-13B Text 13.731 7.541 16.446 80.549 13.884 Image 10.949 6.009 9.583 79.277 10.191 $\ \ \Delta$ Avg. 2.875 1.064 6.091 0.846 4.254 BLIP-2 (Li et al., 2023) Flan-T5-XL Text 20.722 10.511 20.118 85.859 15.827 Image 14.107 6.173 13.376 81.343 11.673 Flan-T5-XXL Text 23.174 12.482 21.891 86.720 19.680 Image 18.398 9.490 16.237 84.190 15.294 $\ \ \Delta$ Avg. 5.695 3.664 6.197 3.522 4.270 GPT-4V (OpenAI, 2023) Text 24.320 10.179 22.748 87.457 10.524 Image 22.675 8.812 20.477 85.495 5.067 $\ \ \Delta$ Avg. 1.644 1.367 2.271 1.961 5.456

### v4

@@ Abstract
Recent advances in instruction-tuned Large Vision-Language Models (LVLMs) have imbued the models with the ability to generate high-level, image-grounded explanations with ease. While such capability is largely attributed to the rich world knowledge contained within the Large Language Models (LLMs), our work reveals their shortcomings in fine-grained visual categorization (FGVC) across six different benchmark settings. Most recent state-of-the-art LVLMs such as LLaVa-1.5, InstructBLIP and GPT-4V not only severely deteriorate in terms of classification performance, e.g., average drop of 65.58 in EM for Stanford Dogs for LLaVA-1.5, but also struggle to generate descriptive visual attributes based on a concept that appears within an input image despite their prominent zero-shot image captioning ability. In-depth analyses show that instruction-tuned LVLMs suffer from modality gap, showing discrepancy when given textual and visual inputs that correspond to the same concept. In an effort to further the community’s endeavor in this direction, we propose a multiple granularity attribute-centric benchmark and training mixture, Finer, which aims to establish a ground to evaluate LVLMs’ fine-grained visual comprehension ability and provide significantly improved explainability.
@@ 1 Introduction
In recent years, Large Vision-Language Models (LVLMs) that are able to generate image-grounded text have seen significant progress. Models such as InstructBLIP (Dai et al., 2023) and LLaVA (Liu et al., 2023b, a) have consistently exhibited strong zero-shot capability in generating image captions, visual reasoning and textual descriptions, and even leveraging external knowledge for complex question answering tasks Marino et al. (2019); Schwenk et al. (2022). Such results across diverse benchmarks indicate that these models, most of them being built on large language models (LLMs) like Vicuna (Chiang et al., 2023), Flan-T5 (Chung et al., 2022), Llama (Touvron et al., 2023), are already equipped with the ability to simultaneously leverage the interplay between the textual parametric knowledge acquired during pre-training and the image understanding ability acquired during instruction-tuning. Notably, all of these models exhibit strong zero-shot task transferability to multiple downstream tasks.
@@
Conventionally, in the computer vision domain, many previous works on fine-grained visual classification (FGVC) (Wei et al., 2022b; Diao et al., 2022; Zhu et al., 2022; Yang et al., 2022; Wang et al., 2022) sought to accurately classify diverse images ranging from different types of birds, plants, animals Van Horn et al. (2015, 2018) and artificial objects such as cars Krause et al. (2013) and aircrafts Maji et al. (2013). In this work, we investigate whether state-of-the-art LVLMs can combine their image understanding ability and rich textual knowledge acquired during pre-training to handle zero-shot FGVC. To our surprise, while the LVLMs perform almost perfectly, e.g., 98.43 for LLaVA-1.5 (13B) on iNaturalist, at superordinate-level granularity (e.g., birds, jets), their classification abilities do not extend to the coarse and finer-grained concepts (e.g., bald eagle, F-22 Raptor), exhibiting substantially deteriorated classification performance (§3.2); 46.91 for coarse-level and 1.56 for fine-level categories on iNaturalist. Our empirical analyses of these models reveal that these models suffer from modality gap. We empirically demonstrate that such discrepancy stems from LVLMs’ limited ability to exploit the rich parametric knowledge given image input, to infer fine-grained concepts. We also show that such constraints lead to diminished fine-grained understanding of the image, preventing these models from generating accurate and detailed visual attributes of the concepts that appear within an image.
@@
We also present an attribute-centric and multiple granularity classification benchmark and training mixture, Finer. Our benchmark constructs concept-indicative attributes for six conventional FGVC benchmarks like iNaturalist Van Horn et al. (2018) and FGVC-Aircrafts Maji et al. (2013) by (i) generating multiple granular concept labels for visual concept recognition, and (ii) constructing a set of visual attributes per fine-grained concept to measure the ability of LVLMs to accurately generate fine-grained concept descriptions given an image. To summarize, our contributions include:
@@
  • We highlight the lack of fine-grained image comprehension ability of instruction-tuned LVLMs across various real-life objects. To the best of our knowledge, we are the first to explore FGVC as an evaluation criteria for these models and their lack of ability thereof.
@@
  • We underscore the persistence of modality gap in state-of-the-art LVLMs by conducting an extensive per-modality-based probing, revealing the discrepancy in how the two modalities are processed by these models (§4).
@@
  • We construct a novel attribute-centric benchmark for FGVC to open up a new direction for future works to measure LVLMs’ modality gap and their granular image understanding capability. Our Finer training mixture and newly proposed prompting technique AttrSeek enable substantially improved zero-shot FGVC performance for GPT-4V (§3.2) and LLaVA-1.5 (§5.3).
@@ 2 Related Work
2.1 Instruction-tuned Large Vision-Language Models
State-of-the-art LVLMs such as LLaVA (Liu et al., 2023a, b), BLIP-2 Li et al. (2023), InstructBLIP (Dai et al., 2023), and closed-source models like GPT-4V OpenAI (2023); Yang et al. (2023) have brought to our attention their zero-shot task solving abilities, especially in downstream tasks such as Visual Question Answering (VQA), reasoning and image captioning, all of which require output generation conditioned on extensive knowledge of the real-world. Based on an intricate interplay between their parametric knowledge and image understanding ability, they are able to generate sensible outputs. However, many of them focus almost exclusively on image captioning and reasoning, most often disregarding the concept recognition tasks traditional computer vision tasks evaluate on.
2.2 Fine-grained Visual Categorization
Previous works approach FGVC with masked image modeling (He et al., 2022; Ryali et al., 2023), concept meta-information injection (Diao et al., 2022), or LLM-generated concept descriptions (Menon and Vondrick, 2023). However, learning of fine-grained visual categories in LVLMs and their ability to elaborate on the fine-grained details of the input image via text generation is yet to be explored. Furthermore, recent works in FGVC have also shown to disregard fine-grained details of images (Krojer et al., 2022) and work poorly on downstream tasks involving localization (Ranasinghe et al., 2023; Zhong et al., 2022) and object attributes (Yuksekgonul et al., 2022).
2.3 Modality Gap in Vision-Language Models
Different from instruction-tuned VLMs like LLaVA and InstructBLIP, contrastively trained VLMs like CLIP Radford et al. (2021) and GLIP (Li et al., 2022) rely on directly minimizing the contrastive objective between visual and textual representations. A recent analytical work (Liang et al., 2022) on CLIP-like models reveals that there is a modality gap between the text and visual modalities, which imposes substantial implications on downstream task performances. Our work shows that such modality gap also exists in LVLMs like LLaVA and InstructBLIP, despite the noticeable architectural difference between models like CLIP and LVLMs discussed in this work.
@@ 2.3 Modality Gap in Vision-Language Models
Figure 3: Fine-grained classification pipeline. At each level, an output from LVLM is injected into the next level prompt. (1) Superordinate-level prompt is used to predict the highest-level category (e.g., bird). (2) Coarse-level prompt is subsequently fed with the predicted output and fed back to the LVLM to generate the next output (e.g., parrot), and (3a) and (4) follow the same steps. (3b) illustrates AttrSeek, a newly proposed prompting scheme in this work, wherein the model is prompted to generate the visual attributes.
@@ 3.1 Evaluation Settings
We assess the accuracy of the generated concept labels given an image and a granularity-specific prompt asking the model to figure out the correct category the concept in the image belongs to. Following previous works on concept classification using auto-regressive models, we employ F1 and Exact Match (EM) scores; note that the EM score used in this work is a modified EM score that parses a sequence of generated text and considers the output label correct if the ground-truth label string exists within a pre-defined maximum number of tokens, $m$ (we set $m=20$).
@@
The models used in this work are as follows: LLaVA-1.5 (Liu et al., 2023b, a), InstructBLIP (Dai et al., 2023), and GPT-4V; the hyperparameter setting for each model is in Appendix A.1. The open-sourced models like LLaVA-1.5 and InstructBLIP follow a generic pipeline of transforming an input image $\mathbf{X_{v}}$ with a frozen vision encoder such as CLIP ViT-L/14 Radford et al. (2021) into an encoded image representation $\mathbf{Z_{v}}$. Then, these models either project $\mathbf{Z_{v}}$ into the language representation space through a learned projection layer $\mathbf{W}$, which becomes $\mathbf{H_{v}}=\mathbf{W}\cdot\mathbf{Z_{v}}$ as in Liu et al. (2023b), or attend over $\mathbf{Z_{v}}$ with learnable queries $\mathbf{Q}$ as in Li et al. (2023); Dai et al. (2023). Such transformed visual representations interact with $\mathbf{X_{instruct}}$, a language instruction, which attends over the image representations (or queries) within the self-attention layers of the LLM component to generate the final output sequence.
@@ 3.2 Brittleness of Vision-Language Models
To evaluate the fine-grained image recognition ability of LVLMs, we measure their classification performance per granularity by prompting the models to generate the correct label for a given concept image as shown in Figure 2. As illustrated in Figure 3, we assess the models’ classification ability across three different granularity levels. In Figure 2, we evidence significant deterioration in terms of classification performance across all the five baselines, with some, e.g., iNaturalist-2021, even reaching near 0% in EM score. While models do perform very well for superordinate-level categories, often achieving 100% in EM, the finer granularity leads to substantially worsened classification performance. In terms of model size, larger models like LLaVA-1.5 (13B) and GPT-4V tend to perform better than smaller models like the 7B variants. For InstructBLIP, the 7B version performs better than the 13B version by a large margin, with the 13B version exhibiting less capable instruction-following ability than the 7B one, potentially due to 7B variant being less prone to overfitting and exhibiting efficiency in simple tasks like classification.
@@
Table 1: Elicitive prompting results (EM;%) on iNaturalist. Prompting techniques like Chain-of-Thought (CoT) on the fine-grained classification (Fine) is unable to improve performance in open-source models.
@@
LLMs like Vicuna (Chiang et al., 2023) and LLama (Touvron et al., 2023) used as textual reasoning components of LVLMs are known to perform better when presented with elicitive prompts like Chain-of-Thought (CoT) (Wei et al., 2022a) that improve model’s reasoning ability. A unifying thought along this line of prompting techniques is to break down a complex problem into a sequence of sub-problems, i.e., divide-and-conquer. Inspired by these prompting techniques, we propose and evaluate our prompting technique for FGVC, AttrSeek. In this simple prompting strategy, we first (i) prompt the models to generate the most distinctive physical attributes visible in the concepts in an image, and (ii) feed the generated set of attributes along with the concept-asking prompt for final prediction as shown in Figure 3. We evaluate the 13B model variants and GPT-4V only on iNaturalist-2021 dataset due to budget constraint of evaluating on the full evaluation set. In Table 1, we do not see much improvement in terms of fine-grained concept classification in the open-source models, with neither CoT nor AttrSeek enhancing the classification performance. However, there is a substantial increase in fine-level performance for GPT-4V when prompted with our simple yet effective AttrSeek scheme. This result suggests that LLaVA-1.5 and InstructBLIP, while they exhibit strong image captioning and reasoning ability, are limited in terms of image-grounded attribute understanding even when provided with elicitive prompts; we further elaborate on the need to finetune the model according to the newly proposed prompting scheme in Section 5.3. This result also suggests that open-source LVLMs may lag behind in instruction-following abilities in comparison to GPT-4V. For additional details on few-shot prompting, see Appendix A.3.
@@ 4 Modality Gap: Discrepancy Between Textual and Visual Modalities
We hypothesize that the lack of zero-shot concept classification ability of LVLMs arises mainly due to the modality gap between textual and visual inputs, preventing the models from leveraging the existing concept-related parametric knowledge when an image of a concept is given. Note that modality gap studied in this work is different from the one previously identified in CLIP-like VLMs (Liang et al., 2022). In this section, we aim to delve into the details of how these models process visual and textual modalities by exploring how well they perform when given only textual descriptions of a concept (§4.1) and how they accurately elaborate what they see in a given image (§4.2). We also perform linear probing (§4.3) against projected and original vision encoder output embeddings to gauge the influence of projection and the subsequent loss of visual information on the modality gap.
Base Model LLM Input Modality Similarity Measurement Metrics ROUGE-1 ROUGE-2 ROUGE-L BertScore AlignScore LLaVA-1.5 (Liu et al., 2023a) Vicuna-7B Text 22.343 11.023 21.161 85.429 18.935 Image 17.468 10.049 17.314 84.499 15.509 Vicuna-13B Text 20.616 10.902 19.647 85.944 15.746 Image 19.661 11.334 19.114 85.704 13.319 $\ \ \Delta$ Avg. 2.920 0.703 2.189 0.585 2.926 InstructBLIP (Dai et al., 2023) Vicuna-7B Text 18.987 9.192 16.577 83.102 15.002 Image 16.018 9.788 11.258 82.681 10.186 Vicuna-13B Text 13.731 7.541 16.446 80.549 13.884 Image 10.949 6.009 9.583 79.277 10.191 $\ \ \Delta$ Avg. 2.875 1.064 6.091 0.846 4.254 GPT-4V (OpenAI, 2023) Text 24.320 10.179 22.748 87.457 10.524 Image 22.675 8.812 20.477 85.495 5.067 $\ \ \Delta$ Avg. 1.644 1.367 2.271 1.961 5.456